OpenStudy (anonymous):

How do you transform the polar equation theta=pi/6 to an equation in rectangular coordinates and identify its shape?

OpenStudy (anonymous):

angle is a contant \(\implies \) it is a line

OpenStudy (anonymous):

@satellite73, Right, but how do you change it into rectangular coordinates? :o

OpenStudy (tkhunny):

Same as you transform any other polar equation. The inverse cosine might give a clue as to which way the line points.

OpenStudy (anonymous):

if i am not mistaken the slope will be \(\tan^{-1}(\frac{\pi}{6})\) @tkhunny sound right?

OpenStudy (tkhunny):

Or the inverse sine (use r = 1) or the inverse tangent. Whatever helps you to figure out that it's a line through the origin and your only task is to find the right angle.

OpenStudy (anonymous):

wouldn't it be tan (pi/6) ?\[\tan (\pi/6) \] ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):